OpenStudy (anonymous):

if 2f(x) - 3f(1/x) = x^2, determine f(2). What would the steps be to solve this?

OpenStudy (anonymous):

You can write two equations one for x=2 one for x=1/2 It will have two variables f(2) and f(1/2) Solving it will give you f(2) and f(1/2)

OpenStudy (anonymous):

I have the answer key and it is saying that f(2) = -7/4

OpenStudy (anonymous):

Might be

OpenStudy (anonymous):

Alright, i'll try it out! thanks!

OpenStudy (anonymous):

If you are stuck let me know and good luck!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yes -7/4 is the correct answer

OpenStudy (anonymous):

So I would have 1. f(2) and 2. f(1/2) and I would subsitute that into the function two times?

OpenStudy (anonymous):

Not really. First set x=2 and write up an equation Than set x=1/2 and do the same.

OpenStudy (anonymous):

After I set x = 2, I get 4 - 3/2 = 4 which doesn't do much for me

OpenStudy (anonymous):

when x=2 2f(2)-3f(1/2)=4

OpenStudy (anonymous):

You do not know what the function f does

OpenStudy (anonymous):

So what would I do after that when I am left with 2f(2)-3f(1/2)=4

OpenStudy (anonymous):

write up the same when x=1/2

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

then you have two equations with two unknown variables and I guess you already learnt how to solve that

OpenStudy (anonymous):

So I can either use the subsitution or elimination method?

OpenStudy (anonymous):

I got it!

OpenStudy (anonymous):

Very good